‏1 kg of Dry saturated steam at 10 bar expand reversibly and polytropically to a pressure of bar if the index of expansion is 1.1 ,
Find the following ‏
1)The value of Entropy at start of the process is ‏
2)The value of Internal Energy at start of the process is ‏
3)The value of Entropy at end of the process is
4) The value of Internal energy at end of the process is
5) The change of entropy between start and end is 6)The value of heat supply and state its supplied or rejected
7)The value of work done is bc
Answer #1

Solution in Given data mand of the steam Finaal 1 kg Initial pressure, Pi = loboro pressure, P 1 bat index of expansion, n= 1Now for Polytropic process we Com) have: 3 P n On V₂ = 4 P2 Dolg444om? V = Vg @lobar V2 = (0.194440x{ let x2= quality (drynesUz can be Up = Sfga S = Sfat x Sf2 = Sp @ 1 hors = 1.3025 KJ Sfg @ 1 bort 6.os 68 KJ Spg? kg-k regk S2 = (1.3025)+ (0.93)(6.0AS = Sy-Si As = 6.935- 6.5864 AS = 0.3486 KJ kg-k (vi) Heat supplied (01-2):- Q,-2 Can be given by: Q-2= (-4₂) + lool Piv,- PQ1-2 [ (2260-2582164) too [(loxo.194440) 12.64 + 171.571 1.1) 01-2=(-222 64) + (367.4) Q1-2 144.76 kJ kJ/kg Ang As 1 <n<r whe(vii) work done (Wire) Wi-2 came be giver by: Wi-2= 100 (PV-P2 V 2) M-) Wi-2=100 100(10x 0.194449) - (1x1.577)] 1.1 - Wi-z=36